OpenStudy (anonymous):

HELPHELPHELP ASAP!! Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (anonymous):

y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20 I dont get where to get 15 or 20 from

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@jhonyy9

jimthompson5910 (jim_thompson5910):

y = mx+b y = -6x+b -3 = -6(-2) + b keep going to solve for b

OpenStudy (anonymous):

y = -6x - 15